if In = 0∫∞ e-xsinn(x)dx , prove that (1+n2)In = n(n-1) In-2 for n >= 2


1
Expert's answer
2022-05-31T12:24:28-0400

"I_n=\int\limits_0^\infty e^{-x}\sin^nx dx
\\
=\int\limits_0^\infty (-e^{-x})^\prime \sin^nx dx
\\=-e^{-x}\sin^nx\big|_0^\infty+\int\limits_0^\infty e^{-x}\cdot n\sin^{n-1}x\cos xdx 
\\
=0+n \int\limits_0^\infty (-e^{-x})^\prime \cdot\sin^{n-1}x\cos xdx 
\\
=-ne^{-x}\sin^{n-1}x\cos xdx \big|_0^\infty+n\int\limits_0^\infty e^{-x}\bigg( (n-1)\sin^{n-2}x\cos^2 x-\sin^n xdx\bigg)dx
\\
=0+n(n-1)\int\limits_0^\infty e^{-x}\sin^{n-2}(1-\sin^2x)dx-n\int\limits_0^\infty e^{-x}\sin^nxdx 
\\
=n(n-1)I_{n-2}-n(n-1)I_n-nI_n"



"I_n=n(n-1)I_{n-2}-n(n-1)I_n-nI_n"



"(n^2+1)I_n=n(n-1)I_{n-2}"